Cullumber, Inc., is expected to grow at a constant rate of 8.00 percent. If the company’s next dividend, which will be paid in a year, is $1.23 and its current stock price is $22.35, what is the required rate of return on this stock?
Answer #1

Values given are

D1 = $1.23 ; Expected dividend

G = 8% ; growth rate

P0 = $22.35 ; current price of stock

R = ?

P0 = D1 / R - G

R = D1 / P0 + G

= 1.23 / 22.35 + 0.08

= 0.055 + 0.08

= 0.135 or 13.5%